The following equation represents the relationship between hours of TV watched per day (x) and the number of situps a person can do (ŷ). ŷ = -1.343 x + 25.146 Use this model to predict the number of situps a person who watches 5.5 hours of TV can do. (to 2 decimal places)
Answer #1

\widehat{y} :  number of situps a person can do

x: hours of TV watched per day

y = -1.343.7 + 25.146

Use this model to predict the number of situps a person who watches 5.5 hours of TV(x=5.5) can do

plugin x =5.5 in the given equation that represents  the relationship between hours of TV watched per day (x) and the number of situps a person can do \widehat{y}

i.e

y = -1.343.+25.146 = -1.343 x 5.5+25.146 = -7.3865 +25.146 = 17.7595

Predicted number of situps a person who watches 5.5 hours of TV can do = 17.76